You have found three investment choices for one year deposit: 9.5% APR compounded monthly, 9.5% APR compounded annually and 8.8% APR compounded daily ((365) days. Compute the EAR for each investment choice.
Answer #1
EAR = ( 1 + r )^n - 1
9.5% APR compounded monthly
EAR= ( 1 + 9.5%/12)^12 - 1
EAR= 9.92%
9.5% APR compounded annually
EAR= ( 1 + 9.5%/1)^1 - 1
EAR= 9.50%
8.8% APR compounded daily
EAR= ( 1 + 8.8%/365)^365 - 1
EAR= 9.20%